Following information is available for Companies Ace Ltd. and Pace Ltd.: (₹ in lacs)

Particulars Ace Ltd. Pace Ltd.

Long term Debt 625 700

Equity 2100 2850

Current assets 450 550

Current liabilities 300 375

Net Profit 115 178

Revenue (net) 355 452

a. Compute Debt-equity ratio, current ratio for both companies.

b. If face value of equity shares of both companies ₹10 each, calculate the Earnings per share ratio for both companies, advising which company is recommended for investment.


1
Expert's answer
2022-03-21T06:21:02-0400

a) Debt Equity Ratio = (Short term Debt + Long Term Debt) / Shareholders Equity 

Ace Ltd 

Debt Equity Ratio = (625+300)/2100

Debt Equity Ratio =925/2100

Debt Equity Ratio = 0.44

Pace Ltd

Debt Equity Ratio = (700+375)/2850

Debt Equity Ratio = 1075/2850

Debt Equity Ratio = 0.38

Current Ratio = Current Assets/Current Liabilities 

Ace Limited Current Ratio = 450/300 Current Ratio = 1.5

Pace Limited  Current Ratio = 550/375 Current Ratio = 1.5

b) Earnings per share (EPS) is a company's net profit divided by the number of common shares it has outstanding.

EPS = Net Income/Weighted Average share outstanding

Ace Ltd

EPS = 115/210 = 0.55

Pace Ltd

EPS = 178/285 = 0.63

I would recommend Pace Ltd Company for investment. Higher earnings per share is always better than a lower ratio because this means the company is more profitable and the company has more profits to distribute to its shareholders.